# Decision-dependent Robust Charging Infrastructure Planning

### v1

@@
Decision-dependent Robust Charging Infrastructure Planning for Light-duty Truck Electrification at Industrial Sites: Scheduling and Abandonment
@@
Many industrial sites rely on diesel-powered light-duty trucks to transport workers and small-scale facilities, which has resulted in a significant amount of greenhouse emissions (GHGs). To address this, we developed a two-stage robust charging infrastructure planning model for electrifying light-duty trucks at industrial sites. The model is formulated as a mixed-integer linear programming (MILP) that optimizes the charging infrastructure, selected from multiple charger types and potential locations, and determines opportunity charging schedules for each truck based on the chosen infrastructure. Given the strict stopping points and schedules at industrial sites, we introduced a scheduling-with-abandonment problem, where trucks forgo charging if their waiting times exceed a maximum threshold. We also further incorporated the impacts of overnight charging and range anxiety on drivers’ waiting and abandonment behaviors. To represent stochastic, heterogeneous parking durations of trucks, we constructed a decision-dependent robust uncertainty set in which parking time variability flexibly depends on drivers’ charging choices. We applied the model in a case study of an open-pit mining site, which plans charger installations in eight zones and schedules a fleet of around 200 trucks. By decomposing the problem into monthly subproblems and using heuristic approaches, for the whole-year dataset, the model achieves an optimality gap of less than 0.1% within a reasonable computation time under diverse uncertainty scenarios.
@@ 1.1 Background
Diesel-powered light-duty trucks are used for transporting and dispatching goods in urban environments and industrial sites. However, their usage could produce a significant amount of GHG emissions. In 2022, light-duty trucks were the most significant GHG emission source of transportation, contributing to 58% of the total emission of the transportation sector in the U.S. \@BBOPcite\@BAP\@BBNU.S. EPA (2024)\@BBCP. Battery electric vehicles (BEVs) play a central role in decarbonizing the transportation sector. Its associated charging infrastructure is currently categorized into two types. The first type is the level 1 and level 2 AC charging (hereafter referred to as ‘slow charging’), which generally requires between four hours and a full day to charge a vehicle battery from empty to 80% state of charge (SoC). The second type is the DC fast charging (hereafter referred to as ‘fast charging’), available at dedicated stations, which requires a relatively complicated installation process. The fast chargers can recharge an empty vehicle battery to 80% SoC in an hour \@BBOPcite\@BAP\@BBNUS Department of Transportation (2025)\@BBCP. Fast charging infrastructure also entails considerable capital investments and installation expenditures, primarily due to the power electronics required. Furthermore, its operational constraints differ from those of slow charging, as charging beyond 80% SoC is strongly discouraged for safety considerations \@BBOPcite\@BAP\@BBNMussa et al. (2017)\@BBCP. Another recharging option is the battery swapping station. The battery swapping station enables EVs to replace depleted batteries with fully charged ones swiftly \@BBOPcite\@BAP\@BBNMak et al. (2013)\@BBCP. However, due to the safety risks associated with battery swapping operations, the feasibility of deploying battery swapping stations at industrial sites should be assessed on a case-by-case basis.
In this paper, we examine the electrification of light-duty trucks at industrial sites for decarbonization. The original diesel-powered trucks are primarily employed for transporting workers and small-scale facilities. The strict operational schedules at industrial sites leave little room for trucks to change any designated parking locations or timetables. As a result, the scheduled trucks are expected to abandon charging if the anticipated waiting time for getting a charger surpasses a specified threshold. This is formulated as a ‘scheduling-with-abandonment’ problem. We conducted a case study of the open-pit mining site with 200 diesel-powered operational trucks. Without modifying any of their GPS data, we extracted their parking and driving patterns based on potential charging points and estimated their energy consumption at half-hour resolutions. Then, we developed a two-stage charging infrastructure planning and scheduling model for the entire truck fleet. The planned EV charging infrastructure satisfies the charging needs of each truck. That is meeting the energy consumption of each truck at any time, given the limited parking durations and battery energy capacity.
@@ 1.2.1 Mixed-type charging infrastructure planning
Existing studies on charging infrastructure planning primarily focus on determining the location, number, and type of EV chargers, as well as the charging schedule using the selected installation. For electric buses and trucks at industrial sites, the charging process primarily involves opportunity or overnight charging, meaning they are only charged during their layovers and overnight parking \@BBOPcite\@BAP\@BBNAbdelwahed et al. (2020)\@BBCP. The vehicle routing is typically not considered in the electric bus scheduling, as buses usually have fixed stop points and timetables. After determining a finite set of candidate locations for potential EV charger deployment based on stop frequency, the number and types of chargers can be optimized according to trip patterns. \@BBOPcite\@BAP\@BBNHu et al. (2022)\@BBCP models a joint optimization problem for locating fast chargers at selected bus stops and optimizing schedules for opportunity charging. The model assumes that buses will be charged during layovers, and the cost takes into account the time-varying electricity price and penalty for additional waiting times. \@BBOPcite\@BAP\@BBNLegault et al. (2025)\@BBCP formulates a multi-period, integrated planning and scheduling problem. The paper models the electric bus timetable as a graph representing their charging and driving pattern, which is considered a deterministic, minimum-cost problem. \@BBOPcite\@BAP\@BBNGkiotsalitis et al. (2025)\@BBCP optimize the EV charger location and number by allocating bus trips to their nearest chargers with available time slots, assuming that buses will be driven to the optimal charger location after the trip ends.
1.2.2 EV charging scheduling with waiting time and abandonment
@@
The scheduling problem is a fundamental and widely studied combinatorial optimization problem, with extensive applications in areas such as supply chain management, service management, and healthcare \@BBOPcite\@BAP\@BBNFonseca et al. (2024), Delarue et al. (2025)\@BBCP. EV charging can be viewed as a service that relies on a reusable resource—namely, the charging infrastructure. The EV charging scheduling problem could thus consider different attributes, such as charger availability, charging time, and waiting times for charging services. For instance, the delay or waiting time can be modeled as a monetary penalty in the objective function \@BBOPcite\@BAP\@BBNHu et al. (2022), Tao et al. (2024), Gkiotsalitis et al. (2025)\@BBCP. When the waiting time exceeds a threshold, the vehicle that committed a time-critical task or trip might abandon charging and leave to find another charging spot, which can be represented as a scheduling problem with abandonment \@BBOPcite\@BAP\@BBNG. Chen et al. (2023)\@BBCP. Another critical aspect of EV charging is range anxiety, referring to drivers’ concern about depleting their battery before reaching a charging station or their destination. Empirical studies suggest that range anxiety may incentivize the installation of additional charging stations or the adoption of larger battery capacities \@BBOPcite\@BAP\@BBNLim et al. (2015)\@BBCP. \@BBOPcite\@BAP\@BBNLu et al. (2025)\@BBCP model the impact of range anxiety by setting a threshold SoC before reaching the destination. \@BBOPcite\@BAP\@BBNTao et al. (2024)\@BBCP formulate the accumulated range anxiety level as a function of SoC depletion along the route and integrates the range anxiety into its objective function.
1.2.3 Decision-dependent robust optimization
@@
In integrated charging infrastructure planning and scheduling problems, several uncertainties arise, including EV arrival and departure times, charging durations, and energy consumption. They can be modeled using stochastic optimization (SO), robust optimization (RO), and distributionally robust optimization (DRO). Among these formulations, RO typically ensures the highest level of reliability, making it particularly suitable for critical infrastructure planning. Different shapes of uncertainty sets, such as box or ellipse, lead to the robust counterparts and tractability \@BBOPcite\@BAP\@BBNBertsimas & Hertog (2022)\@BBCP. \@BBOPcite\@BAP\@BBNHu et al. (2022)\@BBCP model the uncertainty of passengers’ boarding and alighting time during bus trips using box and budget uncertainty sets based on the RO formulation. \@BBOPcite\@BAP\@BBNL. Chen et al. (2024)\@BBCP conduct EV charging scheduling considering the stochastic arrival time of different user groups, where an exponential cone programming formulation is introduced to approximate the upper bound of the stochastic problem. \@BBOPcite\@BAP\@BBNY. Chen & Liu (2023)\@BBCP developed a two-stage stochastic model for planning and operating shared autonomous EV systems, where scenarios represent uncertain user trip requests. An accelerated two-phase Benders decomposition-based algorithm is proposed for solving this two-stage programming problem.
However, these constructed uncertainty sets model uncertainties independent of exogenous factors such as charging decisions. A number of charging installation trials have already shown that charging decisions are very likely to impact parking and waiting durations. For instance, parking duration is strongly correlated with the charging types (e.g., whether it is slow or fast charging) \@BBOPcite\@BAP\@BBNCui et al. (2023), Zhang et al. (2025)\@BBCP. This could be modeled as the decision-dependent uncertainty set, and its uncertainty is endogenous \@BBOPcite\@BAP\@BBNNohadani & Sharma (2018)\@BBCP. \@BBOPcite\@BAP\@BBNNohadani & Sharma (2018)\@BBCP demonstrates that, in general, RO problems with decision-dependent uncertainty are NP-complete and require either global optimization or reformulation techniques such as the big-M method. Similarly, \@BBOPcite\@BAP\@BBNNoyan et al. (2022)\@BBCP introduces a decision-dependent ambiguity set in a machine scheduling problem and examines how different types of uncertainty dependence affect the model formulation. Their findings show that incorporating affine dependency on decision variables can reduce nonlinear constraints through the use of McCormick envelopes and the big-M approach.
Decision Variables User Behavior Uncertainty formulation
@@
Our paper RO
  • Notes: D = Deterministic model; RO = Robust Optimization; SO = Stochastic Optimization.
All the aforementioned work about charging infrastructure planning and scheduling is presented in Table 1.2.3. Compared to these works, our paper makes the following contributions:
  • We developed a two-stage EV charger planning model that optimizes the charger location, quantity, type, and scheduling. The model introduces a scheduling-with-abandonment formulation that reflects the industrial site’s strict operational rules and captures the impacts of overnight parking, range anxiety, and waiting times on the charging scheduling.
  • We developed a decision-dependent uncertainty set to capture the variability in parking durations and their differences across charger types (i.e., slow and fast chargers). Although decision-dependent robust optimization models are typically nonlinear, we demonstrated that the proposed affine dependence on the binary charging decision can be linearized, allowing the problem to be formulated as a MILP.
  • Considering the high computation cost of the scheduling problem over a long optimization horizon, we designed a ‘fix-and-optimize’ heuristic method to forecast parking intervals with high charging demands in the scheduling. Using the representative days or iterative planning by month, the model achieves an optimality gap of less than 0.1% within a reasonable computation time under four uncertainty scenarios.
@@ 2.1 Problem description
The problem is formulated as a two-stage model: the first stage determines charger installations, and the second stage schedules truck charging based on the installed infrastructure. We consider a fleet of electric light-duty trucks $i\in\mathcal{I}$ operating around an industrial site based on the fixed stop points and schedules. As shown in Figure 1, a truck $i$ can be charged during its available half-hour parking time slots $\mathcal{S}_{i,d,t}$ at time $t$ on day $d$ at a charging zone $z$. One truck cannot appear in more than one zone at a time.
Figure 1: The studied multi-zone charging scheduling problem for the light-duty truck fleet; At any half-hour parking time slot $\mathcal{S}$ shown in light green, a truck can either choose to wait, abandon charging, or charge. The effective parking durations pppp are shown in dark green.
Within a parking slot ${S}_{i,d,t}$, a truck may either wait to be charged, abandon charging, or charge directly using one of the available charger types $j\in\mathcal{J}$. In each parking time slot, the effective parking duration is less than half an hour, which restricts the maximum amount of energy that can be charged for this truck during this parking time window. Across the optimization horizon, the total energy of each truck scheduled to be charged must satisfy the truck’s trip-related energy consumption, considering the battery energy balance, battery SoC constraints, and charging power constraints. The traveling distance and energy consumption during parking slots ${S}_{i,d,t}$ is not considered. Furthermore, at any time $t$ on the day $d$, the number of chargers simultaneously used in zone $z$ in type $j$ cannot exceed the number of this type of chargers installed in the zone, $x_{z}^{j}$.
2.2 Set, index, and parameters

### v2

@@
Decision-Dependent Robust Charging Infrastructure Planning and Scheduling for Light-Duty Truck Electrification at Industrial Sites
@@
Many industrial sites and digital logistics platforms rely on diesel-powered light-duty trucks to transport workers and small-scale facilities, which results in a significant amount of greenhouse gas emissions (GHGs). To address this, we develop a robust model for planning charging infrastructure to electrify light-duty trucks at industrial sites. The model is formulated as a mixed-integer linear program (MILP) that optimizes the charging infrastructure selection (across multiple charger types and locations) and determines charging schedules for each truck based on the selected infrastructure. Given the strict stop times and schedules at industrial sites, we introduce a scheduling-with-abandonment problem in which trucks forgo charging if their waiting time exceeds a maximum threshold. We further incorporate the impacts of overnight charging and range anxiety on drivers’ waiting and abandonment behaviors. To model stochastic, heterogeneous parking durations, we classified trucks using machine learning (ML) methods based on contextual and time-location features. We then constructed decision-dependent, feature-driven robust uncertainty sets in which parking-time variability varies flexibly with drivers’ charging choices. These feature-driven sets are applied to two robust optimization formulations with decision-dependent uncertainty (RO-DDU), resulting in distinct outcomes and managerial implications. We conduct a case study at an open-pit mining site to plan charger installations across eight charging zones, serving approximately 200 trucks. By decomposing the problem into a short rolling horizon or using a heuristic approach for the full-year or representative-day dataset, the model achieves an optimality gap of less than 0.1% under diverse uncertainty scenarios.
[…]
Charging infrastructure planning; Decision-dependent uncertainty; ML classification; Contextual (Stochastic) optimization, Scheduling
@@ 1.1 Background
Many industrial sites, such as the mining and metals sectors, and digital logistics platforms, rely on diesel-powered light-duty trucks to transport and dispatch goods. Their use has produced significant GHG emissions. In 2022, light-duty trucks were the largest GHG emission source in the U.S. transportation sector, contributing 58% of total emissions \@BBOPcite\@BAP\@BBNU.S. EPA (2024)\@BBCP. Battery-electric vehicles (BEVs) are a key solution for decarbonization. Currently, the BEV charging infrastructure can be categorized into two types. The first type is Level 1 and Level 2 AC charging (hereafter referred to as “slow charging”), which generally requires between four hours and a full day to charge a vehicle battery from empty to 80% state of charge (SoC). The second type is the direct current fast charging (hereafter referred to as “fast charging”), which has been deployed at certain public stations. Fast chargers can recharge an empty vehicle battery to 80% SoC in about an hour \@BBOPcite\@BAP\@BBNUS Department of Transportation (2025)\@BBCP. However, fast-charging infrastructure entails considerable capital investment and installation expenditures, primarily due to the required power-electronics installations. Fast charging beyond 80% SoC is strongly discouraged for safety reasons \@BBOPcite\@BAP\@BBNMussa et al. (2017)\@BBCP. Another fast recharging option is battery swapping. Battery swapping enables EVs to replace depleted batteries with fully charged ones swiftly \@BBOPcite\@BAP\@BBNMak et al. (2013)\@BBCP. However, compared with adding more trucks to the fleet, deploying battery swapping at industrial sites may yield limited benefits due to operational risks, including potential battery damage during swapping.
In this paper, we develop an integrated planning and scheduling model for charging infrastructure with mixed-type BEV chargers, leveraging contextual stochastic optimization and ML techniques. We comprehensively incorporate heterogeneous BEV driving behaviors in the scheduling, including charging abandonment, overnight charging, and range anxiety. To capture variability in parking duration and its dependence on charging decisions, we construct a feature-based, decision-dependent uncertainty set using ML-based clusters and apply it to two RO-DDU formulations with different assumptions and computational costs. We conduct a case study at an open-pit mining site with 200 operational diesel-powered trucks. Without modifying the GPS data, we extract parking and driving patterns relative to potential charging locations, estimate energy consumption, and efficiently compute the charging infrastructure plan and schedules.
@@ 1.2.1 Mixed-type Charging Infrastructure Planning
Existing studies on charging infrastructure planning primarily focus on determining locations, quantities, and types of BEV chargers, as well as the associated charging schedules given selected installations. For electric buses and trucks at industrial sites, their charging operations typically rely on the opportunity or overnight charging, meaning that vehicles are charged during layovers and overnight parking periods \@BBOPcite\@BAP\@BBNAbdelwahed et al. (2020)\@BBCP. After determining a finite set of locations for EV charger deployment, the numbers and types of chargers can be optimized based on committed trips and available buses. In prior studies, including \@BBOPcite\@BAP\@BBNHu et al. (2022)\@BBCP, \@BBOPcite\@BAP\@BBNLegault et al. (2025)\@BBCP, and \@BBOPcite\@BAP\@BBNGkiotsalitis et al. (2025)\@BBCP, the charging planning objective includes charging infrastructure costs, BEV purchase costs, and operational costs such as scheduling penalties and charging prices. \@BBOPcite\@BAP\@BBNHu et al. (2022)\@BBCP model a joint optimization problem for locating fast chargers at selected bus stops and optimizing schedules during bus layovers, and the objective function incorporates the time-varying electricity price and penalty for the additional waiting times. \@BBOPcite\@BAP\@BBNLegault et al. (2025)\@BBCP formulates a multi-period integrated charging infrastructure planning and scheduling problem for an electric bus fleet. The model accounts for BEV and charger acquisition costs over multiple years, as well as the phased retirement of diesel buses. \@BBOPcite\@BAP\@BBNGkiotsalitis et al. (2025)\@BBCP optimize the mixed-type BEV chargers’ locations and numbers by allocating bus trips to their nearest chargers when they are available, assuming that buses are charged at the depots or the end of their trips.
1.2.2 BEV Charging Scheduling
@@
The classical multi-depot vehicle scheduling problem takes a set of timetabled trips as input and requires that they all be performed by a single vehicle. The BEV charging scheduling problem is considered as a variant of this problem in which battery SoC constraints are imposed. The problem can be modeled as a network with nodes corresponding to discretized battery SoC levels and solved as a shortest path problem \@BBOPcite\@BAP\@BBNDe Vos et al. (2024)\@BBCP. The charging procedure can be viewed as a service scheduling if a fleet of vehicles is involved. The charging problem thus requires consideration of charger availability and waiting times (or queues) for charging services. For instance, the waiting time can be modeled as a monetary penalty in the objective function \@BBOPcite\@BAP\@BBNHu et al. (2022), Tao et al. (2024), Gkiotsalitis et al. (2025)\@BBCP. When the waiting time exceeds a threshold, the vehicle that initiated a time-critical task or trip may abandon charging and seek another charging spot, which can be modeled as abandonment. A unique aspect of BEV charging is range anxiety, the concern among drivers that their battery will deplete before reaching a charging station or their destination. Empirical studies show that range anxiety may incentivize the installation of additional charging stations or the adoption of larger battery capacities \@BBOPcite\@BAP\@BBNLim et al. (2015)\@BBCP. \@BBOPcite\@BAP\@BBNLu et al. (2025)\@BBCP model the impact of range anxiety by setting a threshold SoC before reaching the destination. \@BBOPcite\@BAP\@BBNTao et al. (2024)\@BBCP formulate the accumulated range anxiety level as a function of SoC depletion along the route and integrates this term into the objective function.
1.2.3 Charging Event Characteristics and Uncertainty
@@
Charging events are usually characterized by charging duration and charging power (or energy), and are correlated with diverse features including charging type, hour of day, day of week, location, and vehicle type \@BBOPcite\@BAP\@BBNCui et al. (2023), Zhang et al. (2025)\@BBCP. \@BBOPcite\@BAP\@BBNCui et al. (2023)\@BBCP processed and analyzed more than 220,000 real-world charging records in Beijing. The authors clustered these charging events based on diverse features and structured them as a behavioral tree of fast charging. These clusters are labeled as low-, medium-, and high-priority based on dwell duration. In BEV charging scheduling problems, charging and dwell durations are typically unknown at the time of scheduling. These uncertainties can be modeled using stochastic optimization (SO), robust optimization (RO), and distributionally robust optimization (DRO). Among these formulations, RO typically provides the highest level of reliability, making it particularly suitable for critical infrastructure planning. Different shapes of uncertainty sets, such as boxes or ellipsoids, lead to different robust counterparts and affect tractability \@BBOPcite\@BAP\@BBNBertsimas & Hertog (2022)\@BBCP. \@BBOPcite\@BAP\@BBNHu et al. (2022)\@BBCP model uncertainty in passengers’ boarding and alighting times during bus trips using box and budget uncertainty sets in an RO formulation. \@BBOPcite\@BAP\@BBNL. Chen et al. (2024)\@BBCP consider EV charging scheduling with stochastic arrival times of different EV user groups and introduce an exponential-cone programming formulation to approximate an upper bound. \@BBOPcite\@BAP\@BBNY. Chen & Liu (2023)\@BBCP develop a two-stage stochastic model for planning and operating shared autonomous EV systems and create scenarios for uncertain user trip requests; an accelerated two-phase Benders decomposition-based algorithm is proposed to solve the resulting problem. \@BBOPcite\@BAP\@BBNWang et al. (2023)\@BBCP develop an ML-cluster-based, feature-driven ambiguity set for surgery scheduling using a DRO formulation; the clustering is based on classification and regression trees.
1.2.4 Decision-dependent robust optimization for charging decisions
With electrification, parking durations that previously involved no charging activities are likely to change after charger deployment. Large-scale charging-installation trials conducted in multiple countries have demonstrated that charging decisions significantly affect BEV parking and waiting durations \@BBOPcite\@BAP\@BBNCui et al. (2023), Zhang et al. (2025)\@BBCP. Parking (or charging) duration can thus be modeled using a decision-dependent uncertainty set that depends on endogenous factors, such as charging decisions. \@BBOPcite\@BAP\@BBNNohadani & Sharma (2018)\@BBCP demonstrate that, for simple robust linear problems, RO-DDU is NP-complete and requires either global optimization or reformulation techniques such as the big-M method. \@BBOPcite\@BAP\@BBNLuo & Mehrotra (2020)\@BBCP study decision-dependent DRO models with five types of ambiguity sets, where an instantaneous decision directly impacts right-hand-side uncertainty, including moments and the radius of ambiguity sets. \@BBOPcite\@BAP\@BBNNoyan et al. (2022)\@BBCP propose a decision-dependent ambiguity set for a machine scheduling problem and show that introducing affine dependence on decision variables can substantially reduce the number of constraints induced by McCormick envelopes and big-M reformulations. However, existing studies are largely limited to settings with instantaneous decisions, whereas in practice, uncertainty is influenced by aggregated and expected decision outcomes.
Decision Variables User Behaviors Uncertainty formulation
@@ 1.2.4 Decision-dependent robust optimization for charging decisions
Our paper RO-DDU
  • Notes: D = Deterministic model; RO = Robust Optimization; SO = Stochastic Optimization; RO-DDU = Robust Optimization with Decision-Dependent Uncertainty.
Table 1: Literature Review Summary
All the aforementioned work on charging infrastructure planning and scheduling is summarized in Table 1. Compared with these works, our paper makes the following contributions:
  • We develop an integrated charger planning and scheduling model that optimizes charger locations, quantities, and types, and jointly determines schedules given the selected infrastructure. The model introduces a scheduling-with-abandonment formulation that reflects the industrial site’s strict operational rules and captures the impacts of overnight parking, range anxiety, and waiting times.
  • We develop a feature-based, decision-dependent uncertainty set to capture the variability of charging durations and their dependence on charging decisions. Leveraging ML techniques and contextual information, we restructure time–zone features to preserve the calendar effects of truck trips and group parking durations to efficiently set the bounds of the uncertainty sets.
  • Given the limited discussion in the literature on dependency assumptions in decision-dependent uncertainty problems, we distinguish between two cases in which charging duration depends on either instantaneous charging decisions or cluster-level expected charging decisions. We derive tractable reformulations for both cases and discuss the resulting outcomes and managerial implications.
  • Considering the high computational cost of solving the scheduling problem over a long optimization horizon, we design a “fix-and-optimize” heuristic to forecast parking intervals with potentially high charging demand. Using representative days or a rolling horizon, the model achieves an optimality gap of less than 0.1% within a reasonable computation time under different scenarios.
@@ 2.1 Problem description
In the joint planning and scheduling problem, we consider a fleet of electric light-duty trucks $i\in\mathcal{I}$ operating around an industrial site based on the fixed stop points and schedules during the time period $\mathcal{T}$.
Figure 1: The studied multi-zone charging scheduling problem for the light-duty truck fleet. At any half-hour parking time slot $\mathcal{S}$ shown in light green, A truck can choose to wait, abandon charging, or charge. The effective parking durations $pp$ are shown in dark green.
As shown in Figure 1, a truck $i$ can either wait to be charged, abandon charging, or charge directly using one of the available charger types $j\in\mathcal{J}$ during its available half-hour parking time slots $\mathcal{S}_{i,d,t}$ at time $t$ on day $d$ in charging zone $z$. A truck cannot appear in more than one zone at a time. In each parking time slot, the effective parking duration is less than half an hour, which restricts the maximum amount of energy that can be charged during this window. Across the optimization horizon, the total energy scheduled for each truck must satisfy the truck’s trip-related energy consumption, subject to the battery energy balance, battery SoC constraints, and charging power constraints. Traveling distance and energy consumption during parking slots $\mathcal{S}_{i,d,t}$ are not considered. Furthermore, at any time $t$ on day $d$, the number of chargers of type $j$ used simultaneously in zone $z$ cannot exceed the number of installed chargers of that type in the zone, $x_{z}^{j}$.
2.2 Set, index and parameters